A string is stretched to a length of 337 cm and
both ends are fixed.
If the density of the string is 0.014 g/cm,
and its tension is 1750 N, what is the fundamental frequency?
Answer in units of Hz.
1
Expert's answer
2020-04-23T12:43:11-0400
"f=\frac{1}{2l}\sqrt{\frac{T}{\mu}}"

"f=\frac{1}{2(3.37)}\sqrt{\frac{1750}{0.0014}}=166\ Hz"